# Electromechanical computational model of the human stomach

### v1

@@ 2.4.3 Regional contributions of ICCs and SMCs
The proximal region of the stomach is known to exhibit minimal electrical activity and does not generate peristaltic contractions [2]. To account for heterogeneous contributions of ICCs and SMCs in the normalized form of the monodomain formulation, we introduce dimensionless scaling functions $\chi^{\textup{icc}}$ and $\chi^{\textup{smc}}$, which represents the regional contributions of ICC and SMCs, respectively. For ICCs, which are excitable cells and the primary drivers of slow wave initiation and propagation, $\chi^{\textup{icc}}$ defines the regional capacity to generate electrical signals. It varies smoothly across the domain to reflect known spatial activation patterns. To represent this, we use the same normalized control coordinates $\xi_{l}(\boldsymbol{X})\in[0,1]$ and $\xi_{c}(\boldsymbol{X})\in[0,1]$ (see Eqs. 29 and 30). Consistent with the excitability formulation, $\chi$ is spatially modulated by the longitudinal and circumferential shaping functions $f_{l}(\xi_{l})$ and $f_{c}(\xi_{c})$ (see Eq. 32), allowing a physiologically meaningful heterogeneity between proximal and distal stomach regions. Specifically, it is defined as
@@
The same spatial weighting function is applied to the intracellular coupling term via the field $\chi^{\textup{gap}}(\boldsymbol{X})$, which represents the regional capacity of ICCs to transmit signals to SMCs through gap junctions. Specifically, we set $\chi^{\textup{gap}}(\boldsymbol{X})=\chi^{\textup{icc}}(\boldsymbol{X})$ to account for the reduced density of ICC–SMC gap junctions in the proximal region and the spatial specificity of their electrical coupling [7, 2]. In contrast, for SMCs we assume a uniform distribution of $\chi^{\textup{smc}}(\boldsymbol{X})=1$ throughout the domain, reflecting the assumption of spatially homogeneous SMC properties.
@@ 2.4.4 Electrical diffusivity
The diffusion coefficients $\sigma^{\textup{icc}}$ and $\sigma^{\textup{smc}}$ represent the electrical coupling within ICC and SMC, respectively. Experimental studies report spatial variation in gastric slow wave propagation velocity, with higher velocities observed near the greater curvature and lower velocities along the lesser curvature of the stomach [26, 13]. This observation aligns with anatomical considerations: peristaltic waves propagate along the centerline of the stomach, maintaining a trajectory perpendicular to it. To preserve this orientation, waves must traverse a longer path along the greater curvature than the lesser curvature, necessitating a higher conduction velocity at the greater curvature. Previous studies have attributed the observed propagation patterns and associated conduction velocities primarily to intrinsic frequency gradients within gastric tissue. However, our findings suggest that this explanation is insufficient. Instead, we demonstrate that tissue diffusivity $\sigma^{\textup{icc}}$ plays a more significant role in determining the conduction velocity (cf. D). To capture this effect, we introduce a heterogeneous distribution for $\sigma^{\textup{icc}}$ and $\sigma^{\textup{smc}}$, with increased diffusivity values near the greater curvature—accounting for the higher observed conduction velocities—and decreased diffusivity values near the lesser curvature, consistent with the slower velocities reported in that region.
@@
where $\boldsymbol{X}_{\textup{lesser}}$ denotes a point on the lesser curvature such that $\phi_{\textup{fp}}(\boldsymbol{X}_{\textup{lesser}})=\phi_{\textup{fp}}(\boldsymbol{X})$, i.e., both points lie on the same isocontour of the longitudinal field. The parameters $m_{\sigma}$ and $c_{\sigma}$ are calibrated by fitting conduction velocity and diffusion coefficients using a simplified one-dimensional simulation setup, while experimental conduction velocities are taken from literature. The calibrated parameters are summarized in Table E.7. Combining these relations and solving for $\sigma^{\textup{icc}}(\boldsymbol{X})$ yields the closed‐form expression
@@
In alignment with [16], we set the field of the SMC diffusion coefficient to $\sigma^{\textup{smc}}(\boldsymbol{X})=0.1\sigma^{\textup{icc}}(\boldsymbol{X})$. This reflects the comparatively weaker electrical coupling among SMCs relative to ICCs.
@@ 2.5 Boundary and initial conditions
To model these effects, we apply spring elements acting in the surface-normal direction across the stomach surface with spatially varying stiffness, see Figure 5. Higher stiffness values are applied along the greater and lesser curvatures, to reflect anatomical support to the omentum. Transitions between stiffness regions are smoothed using sigmoid functions based on the harmonic field $\phi_{\textup{gl}}$ (see Section 2.4.2) to avoid stress concentrations and ensure numerical stability.
At anatomical openings, namely the pyloric and lower esophageal sphincter, zero Dirichlet boundary conditions restrict movement to the longitudinal direction, simulating anatomical constraints.
To approximate the mechanical effect of digesta effects in the absence of fluid-structure interaction, we apply spatially varying pressure loads acting normal to the stomach wall. These loads emulate the physiological distension caused by chyme. Higher pressure values are applied in the distal stomach to reflect its role in active mixing and digestion, while lower values are assigned in the proximal stomach for its primary function in accommodation.
@@
where $\boldsymbol{n}$ denotes the outward unit normal. A complete list of the boundary and initial condition values is provided in Tables E.6 and E.5.
@@ 3.1 General
We implemented the coupled gastric motility model in Section 2 into the open-source multiphysics simulation framework 4C [32]. Both the scalar transport problem and the solid mechanics problem are addressed numerically via the finite element method. Linear triangular scalar transport elements and linear quadrilateral nonlinear finite shell elements, based on the $7$-parameter formulation, are used for the discretization. The implementation supports non-conforming meshes at the structure-scalar interface through a dual mortar method [51], enforcing constraints via a Lagrange multiplier approach with node-wise penalty regularization. For further details on the mortar methods, the reader is referred to [51, 52].
@@ 3.2 Convergence study
To verify the correct implementation and assess the numerical robustness of the proposed mathematical model, we conducted systematic convergence studies utilizing QUEENS [53], an open-source Python framework for solver-independent multi-query analyses of large-scale computational models. These studies aimed to determine optimal temporal and spatial discretization parameters for both the electrophysiological and mechanical subproblems, ensuring numerical accuracy and computational efficiency.
@@ 3.2.1 Electrophysiology
It is well established that the conduction velocity $v_{c}$ in reaction-diffusion systems is sensitive to the discretization parameters of the numerical scheme [16], particularly the time step $dt$ and the mesh size $h$. We investigated the convergence behavior of $v_{c}$ in a 11D gastric electrophysiology model. Mechanical coupling was excluded from this study to isolate electrophysiological properties. The computational domain spans the length of the stomach from the pacemaker region to the pylorus and is defined as a 11D line with $x\in[$0.0\text{\,}\frac{\mathrm{mm}}{}$,$250.0\text{\,}\frac{\mathrm{mm}}{}$\$]. No-flux boundary conditions are applied at both ends to reflect the electrically quiescent nature of the fundus and the pylorus. The intrinsic frequencies of ICCs are prescribed using a Gaussian-like function, as described in Section 2.4.2, with parameters listed in Table E.1.
@@
To avoid boundary effects, all calculations were performed after $450\text{\,}\frac{\mathrm{s}}{}$ of simulated time, ensuring a fully entrained slow wave pattern. The results, summarized in Figure 6a, illustrate the convergence behavior of $v_{c}$ with decreasing time step and mesh size. The black contour line indicates a $5\%$ relative deviation from the reference value obtained with the finest spatio-temporal resolution in this analysis. Based on these results, we selected $dt=$ $0.1\text{\,}\frac{\mathrm{s}}{}$ and $h=$ $0.5\text{\,}\frac{\mathrm{mm}}{}$ to solve the electrophysiological problem in all subsequent examples. This choice provides a compromise between numerical accuracy and computational cost, maintaining relative errors below $5\%$. This analysis confirms that the selected time step and the mesh size yield conduction velocities within physiologically plausible ranges [26].
@@ 3.2.2 Coupled electromechanics
To determine appropriate numerical parameters for the solid mechanics component of the coupled electromechanical model, we performed a second convergence analysis using a coupled gastric motility simulation. Building on the electrophysiological convergence study, we next assessed the mechanical response under coupled conditions. In contrast to the previous study, this analysis was carried out on a 22D cylindrical wedge, as defined in Section 3.3.2, to capture physiologically relevant fiber orientations. The mechanical response was modeled using bilinear quadrilateral $7$-parameter shell elements. The constitutive framework follows the finite elasticity formulation presented in Section 2, incorporating a constrained mixture model with active strain and prestress. All model parameters were adopted from Table E.2, which summarizes the baseline values of the constrained mixture model. The parameters that were specifically modified here are listed in Table E.3. In particular, identical material properties were applied in both the circumferential and longitudinal directions, using those defined for the longitudinal direction.
@@ 3.3.1 Membrane - solid - shell
The choice of finite element formulation significantly impacts accuracy, computational efficiency, and ease of implementation. While 33D solid elements capture detailed representations of mechanical behavior, they are prone to locking phenomena and pose challenges in meshing thin geometries. Ensuring a proper aspect ratio often requires a fine in-plane mesh, leading to increased degrees of freedom and higher computational costs. Moreover, solid meshes tend to become distorted under large deformations, potentially causing inaccurate or unstable simulations; in the worst case, this necessitates re-meshing to maintain solution fidelity [54]. To address these issues, 22D manifold representations—membrane and shell elements—can be adopted. However, their comparative performance, particularly in gastrointestinal biomechanics, remains underexplored. This section evaluates three element formulations for gastric electromechanics: (i) a nonlinear finite membrane element model [55], (ii) a 33D solid eight-noded hexahedral (HEX88) elements using the F-Bar method to avoid volumetric locking in the case of near incompressibility, and (iii) the $7$-parameter shell element model introduced above. To mitigate locking pathologies (such as in-plane shear, transverse shear, membrane, Poisson thickness, and curvature thickness locking) associated with the finite shell element implementation, we employ both the EAS [56, 35, 34] and the assumed natural strain (ANS) [57] methods. The original $7$-parameter shell formulation can produce ill-conditioned stiffness matrices due to the thin-walled nature and large eigenvalue spectrum. To address this, we incorporate the scaled director conditioning approach by [58], a mechanically motivated preconditioner tailored for thin-walled structures discretized with continuum-based elements.
@@ 3.3.2 Setup
The computational domain is a cylindrical wedge with an opening angle of $0.563^{\circ}$. The cylinder’s axis aligns with the $x$-axis. It measures $250\text{\,}\frac{\mathrm{mm}}{}$ in length, has a radius of $50.930\text{\,}\frac{\mathrm{mm}}{}$, and a wall thickness of $3.5\text{\,}\frac{\mathrm{mm}}{}$. The computational meshes are generated using Coreform Cubit [59].
@@
A unidirectional excitability distribution is applied (see Table E.1), neglecting circumferential propagation to focus on the fully developed contraction wave state. No-flux boundary conditions are applied for the electrophysiology model at both ends of the cylinder. For the solid mechanics problem, zero Dirichlet boundary conditions are imposed axially at both ends, permitting only radial displacements.
@@
Here, $\Psi_{\textup{2D}}$ is used for simulations with incompressible membrane elements, while $\Psi_{\textup{3D}}$ corresponds to the coupled Neo-Hookean formulation (see Section 2.2.3), and is applied to simulations involving $7$-parameter shell elements or solid HEX88 elements. Material constants are taken from Eq. 43 and Table E.4.
@@ 3.3.3 Results
Figure 7b compares the radial displacements $d_{r}$ along the axial position $x$ for different element formulations. Displacements from the HEX88 model are extracted at the midsurface to enable comparison with 22D formulations. The $7$-parameter shell and the HEX88 elements yield nearly identical displacement profiles, while membrane elements predict increased displacement magnitudes and exhibit notable thickness changes due to their inability to resist bending. These discrepancies stem from the kinematically enforced incompressibility in membrane formulations, which does not adequately capture out-of-plane bending modes. The changes of the thickness further illustrate the limitations of membrane elements in reproducing complex 33D deformation states (cf. Figure 7c). In contrast, the shell elements capture bending effects, providing closer agreement with HEX88 elements, although slight differences persist due to the homogeneous thickness assumptions and locking treatment. To quantify computational efficiency, simulation runtimes were compared: membrane elements required $6890\text{\,}\frac{\mathrm{s}}{}$, shell elements $37\,970\text{\,}\frac{\mathrm{s}}{}$, and solid elements $172\,600\text{\,}\frac{\mathrm{s}}{}$. These results highlight the trade-off between accuracy and computational cost. While solid elements yield the most accurate kinematic representation, the associated computational burden makes them impractical for large-scale or patient-specific simulations. Conversely, membrane elements offer substantial performance improvements but are susceptible to geometric instabilities such as buckling, particularly under higher contraction strengths or complex boundary conditions. The shell elements demonstrate the most favorable compromise, combining geometric accuracy, bending and membrane capability, and numerical robustness with manageable computational cost. This balance makes the shell formulation particularly suitable for simulations of gastric motility, where thin-walled morphologies, nonlinear deformations, and large domains must be resolved efficiently.

### v2

@@ 2.4.3 Regional contributions of ICCs and SMCs
The proximal region of the stomach is known to exhibit minimal electrical activity and does not generate peristaltic contractions [Grady-2021-GastricConductionReview]. To account for heterogeneous contributions of ICCs and SMCs in the normalized form of the monodomain formulation, we introduce dimensionless scaling functions $\chi^{\textup{icc}}$ and $\chi^{\textup{smc}}$, which represents the regional contributions of ICC and SMCs, respectively. For ICCs, which are excitable cells and the primary drivers of slow wave initiation and propagation, $\chi^{\textup{icc}}$ defines the regional capacity to generate electrical signals. It varies smoothly across the domain to reflect known spatial activation patterns. To represent this, we use the same normalized control coordinates $\xi_{l}(\boldsymbol{X})\in[0,1]$ and $\xi_{c}(\boldsymbol{X})\in[0,1]$ (see Eqs. 29 and 30). Consistent with the excitability formulation, $\chi$ is spatially modulated by the longitudinal and circumferential shaping functions $f_{l}(\xi_{l})$ and $f_{c}(\xi_{c})$ (see Eq. 32), allowing a physiologically meaningful heterogeneity between proximal and distal stomach regions. Specifically, it is defined as
@@
The same spatial weighting function is applied to the intracellular coupling term via the field $\chi^{\textup{gap}}(\boldsymbol{X})$, which represents the regional capacity of ICCs to transmit signals to SMCs through gap junctions. Specifically, we set $\chi^{\textup{gap}}(\boldsymbol{X})=\chi^{\textup{icc}}(\boldsymbol{X})$ to account for the reduced density of ICC–SMC gap junctions in the proximal region and the spatial specificity of their electrical coupling [Sanders-2016-RegulationSMCFunction, Grady-2021-GastricConductionReview]. In contrast, for SMCs we assume a uniform distribution of $\chi^{\textup{smc}}(\boldsymbol{X})=1$ throughout the domain, reflecting the assumption of spatially homogeneous SMC properties.
@@ 2.4.4 Electrical diffusivity
The diffusion coefficients $\sigma^{\textup{icc}}$ and $\sigma^{\textup{smc}}$ represent the electrical coupling within ICC and SMC, respectively. Experimental studies report spatial variation in gastric slow wave propagation velocity, with higher velocities observed near the greater curvature and lower velocities along the lesser curvature of the stomach [EgbujiGrady-2010-OriginPropagationSW, Hosseini2023]. This observation aligns with anatomical considerations: peristaltic waves propagate along the centerline of the stomach, maintaining a trajectory perpendicular to it. To preserve this orientation, waves must traverse a longer path along the greater curvature than the lesser curvature, necessitating a higher conduction velocity at the greater curvature. Previous studies have attributed the observed propagation patterns and associated conduction velocities primarily to intrinsic frequency gradients within gastric tissue. However, our findings suggest that this explanation is insufficient. Instead, we demonstrate that tissue diffusivity $\sigma^{\textup{icc}}$ plays a more significant role in determining the conduction velocity (cf. E). To capture this effect, we introduce a heterogeneous distribution for $\sigma^{\textup{icc}}$ and $\sigma^{\textup{smc}}$, with increased diffusivity values near the greater curvature—accounting for the higher observed conduction velocities—and decreased diffusivity values near the lesser curvature, consistent with the slower velocities reported in that region.
@@
where $\boldsymbol{X}_{\textup{lesser}}$ denotes a point on the lesser curvature such that $\phi_{\textup{fp}}(\boldsymbol{X}_{\textup{lesser}})=\phi_{\textup{fp}}(\boldsymbol{X})$, i.e., both points lie on the same isocontour of the longitudinal field. The parameters $m_{\sigma}$ and $c_{\sigma}$ are calibrated by fitting conduction velocity and diffusion coefficients using a simplified one-dimensional simulation setup, while experimental conduction velocities are taken from literature. The calibrated parameters are summarized in Table F.7. Combining these relations and solving for $\sigma^{\textup{icc}}(\boldsymbol{X})$ yields the closed‐form expression
@@
In alignment with [brandstaeter2018a], we set the field of the SMC diffusion coefficient to $\sigma^{\textup{smc}}(\boldsymbol{X})=0.1\sigma^{\textup{icc}}(\boldsymbol{X})$. This reflects the comparatively weaker electrical coupling among SMCs relative to ICCs.
@@ 2.5 Boundary and initial conditions
To model these effects, we apply spring elements acting in the surface-normal direction across the stomach surface with spatially varying stiffness, see Figure 6. Higher stiffness values are applied along the greater and lesser curvatures, as well as close to the pyloric sphincter, to reflect anatomical support to the omentum and the relatively stiff pylorus tissue. Transitions between stiffness regions are smoothed using sigmoid functions based on the harmonic field $\phi_{\textup{gl}}$ (see Section 2.4.2) to avoid stress concentrations and ensure numerical stability.
At anatomical openings, namely the pyloric and lower esophageal sphincter, zero Dirichlet boundary conditions are imposed only in the longitudinal direction, restricting axial displacement while allowing radial and circumferential deformation. This configuration reflects the anatomical support provided by the adjoining duodenum and esophagus, which constrains motion along the longitudinal axis while permitting local wall deformation during peristalsis. In addition, spring-type boundary conditions are applied in the surface-normal direction to represent the elastic compliance of the pyloric sphincter region, preventing excessive displacement and ensuring controlled radial contraction. This combined approach reproduces the limited longitudinal motion observed in MRI studies [Hosseini2023] and preserves physiologically realistic deformation of the distal stomach.
To approximate the mechanical effect of digesta effects in the absence of fluid-structure interaction, we apply spatially varying pressure loads acting normal to the stomach wall. These loads emulate the physiological distension caused by chyme. Higher pressure values are applied in the distal stomach ($25\text{\,}\frac{\mathrm{mmHg}}{}$) to reflect its role in active mixing and digestion, while lower values are assigned in the proximal stomach ($10\text{\,}\frac{\mathrm{mmHg}}{}$) for its primary function in accommodation. These values lie within the physiological intragastric pressure range [Janssen-2011-IntragastricPressure, Wali-2013-PatientFullStomach] and, together with our prestressing procedure, ensure a realistic initial stress state of the gastric wall. This approach represents a first-step approximation of digesta effects, allowing us to focus on the electromechanical behavior. It still captures, to a certain degree, the essential mechanical influence of the chyme, thereby enabling stable and physiologically plausible electromechanical simulations.
@@
where $\boldsymbol{n}$ denotes the outward unit normal. A complete list of the boundary and initial condition values is provided in Tables F.6 and F.5.
@@ 3.1 General
We implemented the coupled gastric motility model in Section 2 into the open-source multiphysics simulation framework 4C [4C]. Both the scalar transport problem and the solid mechanics problem are addressed numerically via the finite element method. Linear triangular scalar transport elements and linear quadrilateral nonlinear finite shell elements, based on the $7$-parameter formulation, are used for the discretization. The implementation supports non-conforming meshes at the structure-scalar interface through a dual mortar method [Popp-2009-Mortar], enforcing constraints via a Lagrange multiplier approach with node-wise penalty regularization. This approach ensures stable and accurate transfer of information between electrophysiology and mechanics meshes, while allowing independent mesh refinement. In particular, since the simulation cost is largely driven by the solid mechanical problem, a coarser mesh can be used for this discretizations, improving computational efficiency without compromising solution fidelity. For further details on the mortar methods, the reader is referred to [Popp-2009-Mortar, Popp-2010-DualMortar].
@@ 3.2 Convergence study
To verify the correct implementation and assess the numerical robustness of the proposed mathematical model, we conducted systematic convergence studies utilizing QUEENS [Biehler2025], an open-source Python framework for solver-independent multi-query analyses of large-scale computational models. These studies aimed to determine optimal temporal and spatial discretization parameters for both the electrophysiological and mechanical subproblems, ensuring numerical accuracy and computational efficiency.
@@ 3.2.1 Electrophysiology
It is well established that the conduction velocity $v_{c}$ in reaction-diffusion systems is sensitive to the discretization parameters of the numerical scheme [brandstaeter2018a], particularly the time step $dt$ and the mesh size $h$. We investigated the convergence behavior of $v_{c}$ in a 11D gastric electrophysiology model. Mechanical coupling was excluded from this study to isolate electrophysiological properties. The computational domain spans the length of the stomach from the pacemaker region to the pylorus and is defined as a 11D line with $x\in[$0.0\text{\,}\frac{\mathrm{mm}}{}$,$250.0\text{\,}\frac{\mathrm{mm}}{}$\$]. No-flux boundary conditions are applied at both ends to reflect the electrically quiescent nature of the fundus and the pylorus. The intrinsic frequencies of ICCs are prescribed using a Gaussian-like function, as described in Section 2.4.2, with parameters listed in Table F.1.
@@
To avoid boundary effects, all calculations were performed after $450\text{\,}\frac{\mathrm{s}}{}$ of simulated time, ensuring a fully entrained slow wave pattern. The results, summarized in Figure 7a, illustrate the convergence behavior of $v_{c}$ with decreasing time step and mesh size. The black contour line indicates a $5\%$ relative deviation from the reference value obtained with the finest spatio-temporal resolution in this analysis. Based on these results, we selected $dt=$ $0.1\text{\,}\frac{\mathrm{s}}{}$ and $h=$ $0.5\text{\,}\frac{\mathrm{mm}}{}$ to solve the electrophysiological problem in all subsequent examples. This choice provides a compromise between numerical accuracy and computational cost, maintaining relative errors below $5\%$. This analysis confirms that the selected time step and the mesh size yield conduction velocities within physiologically plausible ranges [EgbujiGrady-2010-OriginPropagationSW].
@@ 3.2.2 Coupled electromechanics
To determine appropriate numerical parameters for the solid mechanics component of the coupled electromechanical model, we performed a second convergence analysis using a coupled gastric motility simulation. Building on the electrophysiological convergence study, we next assessed the mechanical response under coupled conditions. In contrast to the previous study, this analysis was carried out on a 22D cylindrical wedge, as defined in Section 3.3.2, to capture physiologically relevant fiber orientations. The mechanical response was modeled using bilinear quadrilateral $7$-parameter shell elements. The constitutive framework follows the finite elasticity formulation presented in Section 2, incorporating a constrained mixture model with active strain and prestress. All model parameters were adopted from Table F.2, which summarizes the baseline values of the constrained mixture model. The parameters that were specifically modified here are listed in Table F.3. In particular, identical material properties were applied in both the circumferential and longitudinal directions, using those defined for the longitudinal direction.
@@ 3.3.1 Membrane - solid - shell
The choice of finite element formulation significantly impacts accuracy, computational efficiency, and ease of implementation. While 33D solid elements capture detailed representations of mechanical behavior, they are prone to locking phenomena and pose challenges in meshing thin geometries. Ensuring a proper aspect ratio often requires a fine in-plane mesh, leading to increased degrees of freedom and higher computational costs. Moreover, solid meshes tend to become distorted under large deformations, potentially causing inaccurate or unstable simulations; in the worst case, this necessitates re-meshing to maintain solution fidelity [Liu2024]. To address these issues, 22D manifold representations—membrane and shell elements—can be adopted. However, their comparative performance, particularly in gastrointestinal biomechanics, remains underexplored. This section evaluates three element formulations for gastric electromechanics: (i) a nonlinear finite membrane element model [Gruttmann-1992-membrane], (ii) a 33D solid eight-noded hexahedral (HEX88) elements using the F-Bar method to avoid volumetric locking in the case of near incompressibility, and (iii) the $7$-parameter shell element model introduced above. To mitigate locking pathologies (such as in-plane shear, transverse shear, membrane, Poisson thickness, and curvature thickness locking) associated with the finite shell element implementation, we employ both the EAS [AndelfingerRamm-1993-EAS, Büchter-1994-3DExtensionEAS, SimoRifai-1990-EAS] and the assumed natural strain (ANS) [BatheDvorkin-1985-ANS] methods. The original $7$-parameter shell formulation can produce ill-conditioned stiffness matrices due to the thin-walled nature and large eigenvalue spectrum. To address this, we incorporate the scaled director conditioning approach by [Gee-2005-SDC], a mechanically motivated preconditioner tailored for thin-walled structures discretized with continuum-based elements.
@@ 3.3.2 Setup
The computational domain is a cylindrical wedge with an opening angle of $0.563^{\circ}$. The cylinder’s axis aligns with the $x$-axis. It measures $250\text{\,}\frac{\mathrm{mm}}{}$ in length, has a radius of $50.930\text{\,}\frac{\mathrm{mm}}{}$, and a wall thickness of $3.5\text{\,}\frac{\mathrm{mm}}{}$. The computational meshes are generated using Coreform Cubit [coreform2023].
@@
A unidirectional excitability distribution is applied (see Table F.1), neglecting circumferential propagation to focus on the fully developed contraction wave state. No-flux boundary conditions are applied for the electrophysiology model at both ends of the cylinder. For the solid mechanics problem, zero Dirichlet boundary conditions are imposed axially at both ends, permitting only radial displacements.
@@
Here, $\Psi_{\textup{2D}}$ is used for simulations with incompressible membrane elements, while $\Psi_{\textup{3D}}$ corresponds to the coupled Neo-Hookean formulation (see Section 2.2.3), and is applied to simulations involving $7$-parameter shell elements or solid HEX88 elements. Material constants are taken from Eq. 43 and Table F.4.
@@ 3.3.3 Results
Figure 8b compares the radial displacements $d_{r}$ along the axial position $x$ for different element formulations. Displacements from the HEX88 model are extracted at the midsurface to enable comparison with 22D formulations. The $7$-parameter shell and the HEX88 elements yield nearly identical displacement profiles, while membrane elements predict increased displacement magnitudes and exhibit notable thickness changes due to their inability to resist bending. These discrepancies stem from the kinematically enforced incompressibility in membrane formulations, which does not adequately capture out-of-plane bending modes. The changes of the thickness further illustrate the limitations of membrane elements in reproducing complex 33D deformation states (cf. Figure 8c). In contrast, the shell elements capture bending effects, providing closer agreement with HEX88 elements, although slight differences persist due to the homogeneous thickness assumptions and locking treatment. Table 1 summarizes both the accuracy and computational efficiency of the different element formulations. Maximum absolute errors of radial displacement and thickness relative to the HEX88 reference are reported, with spatial error distributions consistent with those shown in Figure D.1. Relative runtimes and the number of degrees of freedom (DOFs) are also listed. The shell formulation achieves comparable accuracy with roughly three times fewer DOFs and only 2222% of the HEX88 runtime, representing an appealing compromise between accuracy and efficiency. We note that the HEX88 solution is employed as a high-fidelity numerical reference rather than an exact analytical solution. Our convergence study indicates that further mesh refinement produces only marginal changes in the results, confirming that the HEX88 mesh provides a sufficiently accurate reference for the comparisons presented here. Solid elements nevertheless yield the most accurate kinematic representation, but their associated computational burden makes them impractical for large-scale or patient-specific simulations. Conversely, membrane elements offer substantial performance improvements but are susceptible to geometric instabilities such as buckling, particularly under higher contraction strengths or complex boundary conditions. The shell formulation offers an advantageous compromise, combining geometric accuracy, bending and membrane capability, and numerical robustness with manageable computational cost. This balance makes the shell formulation particularly suitable for simulations of gastric motility, where thin-walled morphologies, nonlinear deformations, and large domains must be resolved efficiently.
Table 1: Comparison of different element formulations: degrees of freedom (DOFs), relative computational runtime, and maximum absolute errors of radial displacement $d_{r}$ and wall thickness $h_{t}$ along the longitudinal axis, relative to HEX88.
Element formulation Relative DOFs Relative runtime $\max|d_{r}-d_{r}^{\text{HEX8}}|$ $\max|h_{t}-h_{t}^{\text{HEX8}}|$
[%] [%] [$m$] [$m$]
HEX88 (reference) $100$ $100$ $0.0$ $0.0$
Membrane $14.2$ 44 $0.713$ $1.063$
Shell $7$-p $28.5$ $22$ $0.561$ $0.657$